OpenStudy (anonymous):

The sum of a number and two times a smaller number is 62. Three times the bigger number exceeds the smaller number by 116

OpenStudy (anonymous):

The bigger number is ? The smaller number is ?

OpenStudy (kenshin):

let the first number = x, the smaller number = y 1) x+2y=62 2) 3x-y=116 add (1) to 2x(2) you get 7x=232+62=294, x=42, then backsubstitute x into 1) and you find 42+2y=62, 2y=20, y=10 so the bigger number is 42 the smaller number is 10

OpenStudy (anonymous):

Thanks bro